# Generalised Fermat equation: a survey of solved cases

### v1

@@ 4.2 The case $a=c=1$
In this case, equation (4.1) is of the form
@@
The equation (4.21) with $r=2$, $b=2^{\alpha}$ for $\alpha\geq 2$, and $p\geq 7$ prime was studied by Cohen [45, Section 15.3] who obtained the following result.
@@
Theorem 4.35 describes pairwise coprime solutions to (4.22) under the stated conditions, but this equation can have primitive solutions that are not pairwise coprime, e.g. $(x,y,z)=(2,1,3\cdot 2^{(p-3)/2})$ for $\alpha=p-3$, see [81]. In 2004, Ivorra [82, Chapter 1] proves that there are no other such examples, which completely solves equation (4.22) for $\alpha\geq 2$ and $p\geq 7$. Siksek also independently proved the same result in [126].
@@
If $1\leq\alpha\leq 4$ then for the equation (4.22) with $p=5$, that is,
@@
Bennett, Vatsal and Yazdani [14] considered the equation (4.21) with $r=3$ and $b$ being a prime power, that is,
$x^{p}+s^{\alpha}y^{p}=z^{3}$ (4.23)
@@
If $s$ and $p$ are primes such that $s\neq s_{1}^{3}\pm 3^{s_{2}}$ for any integers $s_{1}$ and $s_{2}$ with $s_{2}\neq 1$, $p>s^{2s}$ and $\alpha\geq 0$, then the equation (4.23) has no solutions in coprime integers $x$, $y$ and $z$ with $|xy|>1$.
@@
Let $\alpha\geq 1$ be an integer, $s\in\{5,11,13,23,29,31,41,43,47,53,59,61,67,71,79,83,97\}$ and let $p\geq 11$ be a prime not dividing $s^{2}-1$. Assume also that
$(s,p)\not\in\{(13,19),(29,11),(43,13),(47,13),(59,11),(61,61),(67,73),(79,97),(97,13),(97,79)\}.$
Then equation (4.23) has no solutions in coprime integers $x$, $y$ and $z$ with $|xy|>1$.
So far in this section we considered the cases $r=2$ and $r=3$. For the next case $r=5$, we have the following result [132, Theorem 2.8].
@@ 4.3 The case $a=1$
In this instance, equation (4.1) reduces to
@@
We first consider the case $r=2$, which reduces equation (4.24) to
@@
Ivorra [81] studies equation (4.25) with $b=2^{\beta}$ for $10\leq\beta\leq p-$ 1, $c=2$ and $p\geq 7$ prime, that is,
@@
For prime $p\geq 7$ and integer $11\leq\beta\leq p-$ 1, the equation (4.26) has no non-trivial primitive solutions.
Let prime $p\geq 7$ and $\alpha\geq 2$. The task of finding all primitive integer solutions to the equation (4.25) with $b=2^{\alpha}$ and $c=3$, that is,
@@
is left as an exercise to the reader in [45, Section 15.8]. The authors of [149] prove that the equation has no solutions in coprime non-zero integers $x,y,z$ with $|xy|>1$.
Bennett and Skinner [13] obtained the following results concerning equation (4.25).
@@
$\alpha\geq\alpha_{0}$, $p>c$ and $(c,\alpha,p)\neq(11,3,13)$, then equation (4.25) has no solutions in non-zero pairwise coprime integers $(x,y,z)$ with $xy\neq\pm 1$.
@@
Suppose that $p\geq 11$ is prime and $\alpha$ is a non-negative integer. If $b\in\{5^{\alpha},11^{\alpha},13^{\alpha}\}$ and $p$ is not a divisor of $b$, then equation (4.25) with $c=2$ has no solution in non-zero pairwise coprime integers $(x,y,z)$.
The equation (4.25) with $b$ a power of $2$ and $c$ a prime was studied by Zhang [146], who proved the following theorem.
@@
Bennett et.al [14] studied (4.28) with $b=s^{\alpha}$ and $c=3^{\beta}$ for prime $s$ and positive integers $\alpha,\beta$, that is,
@@
If $s$ and $p$ are primes such that $s\not\in\{5,3t^{3}\pm 1,9t^{3}\pm 1:t\in\mathbb{N}\}$, $p>s^{28s}$ and $\alpha,\beta$ are positive integers with $3\nmid\beta$, then the equation (4.29) has no solutions in coprime integers $x$, $y$ and $z$ with $|xy|>1$.
@@
If $p\geq 7$ is prime, $s\in\{7,11,13\}$ with $(s,p)\not\in\{(7,13),(13,7)\}$ and $\alpha,\beta$ are positive integers with $3\nmid\beta$, then the equation (4.29) has no solutions in coprime integers $x$, $y$ and $z$ with $|xy|>1$.
Bennett et.al [14] also studied equation
@@
and $\alpha$ and $\beta$ are non-negative integers, then the equation (4.30) has no solutions in coprime integers $x$, $y$ and $z$ with $|xy|>1$, unless $(|x|,|y|,\alpha,p,|s^{\beta}z^{3}|)=$ $(2,1,1,7,125)$.
Krawciów [92] studied equation (4.28) with $b=s^{\alpha}$, that is,
@@
and solved it under some conditions on the parameters. To state the conditions, we must first introduce some notation. Let $s_{1},\dots,s_{k}$ be odd primes, such that $3\in\{s_{1},\dots,s_{k}\}$. Then let $P(s_{1},\dots,s_{k})$ denote the set of primes $x$ satisfying any of the $2^{k+1}$ equations
$x^{\alpha}-{s_{1}}^{\alpha_{1}}\cdots{s_{k}}^{\alpha_{k}}y^{3}=\pm 1,\quad(1\leq\alpha_{i}\leq 2,\,\,i=1,\dots,k).$
for some integers $\alpha>1$ and $y>1$. Theorem 4.34 implies that the set $P(s_{1},\dots,s_{k})$ is always finite. For example, $P(3)=\{2,5\}$, see [92].
@@
Let $c=\prod_{i=1}^{k}s_{i}^{\gamma_{i}}$ be a prime factorization of a positive cube-free odd integer $c$ which is divisible by $3$, $\alpha$ be a positive integer, $p$ is a prime, and let $C=\mathrm{rad}(c)$. If $s$ is a prime such that $s\not\in P(s_{1},\dots,s_{k})$ and $s\neq\prod_{i=1}^{k}s_{i}^{\alpha_{i}}t^{3}\pm 1$ $(1\leq\alpha_{i}\leq 2,i=1,\dots,k)$ for any integer $t$, and if $p>(sC)^{10sC^{2}}$, then equation (4.31) has no non-trivial primitive integer solutions.
@@ 4.4 The case $c=1$
In this instance, equation (4.1) reduces to
@@
4.4.1 Equation (4.32) with fixed $p$
The equation (4.32) with $p=3$ was considered by Bennett and Dahmen which led to the following partial result [9, Theorem 13.2].
@@
has no solution in non-zero integers $x$, $y$ and $z$ (with $\gcd(x,y)=1$) and prime $r\equiv 1$ (mod 3) with $r\geq r_{0}$. If all solutions to (4.33) with $x$ and $y$ coprime integers have $\alpha_{2}\leq 4$, then there exists an effectively computable constant $r_{1}$ such that equation (4.34) has no solutions in odd integers $x$, $y$ and prime $r\geq r_{1}$.
@@
4.4.2 Equation (4.32) with fixed $r$
In 2004, Bennett and Skinner [13] considered equation (4.32) with $r=2$, and $p\geq 11$ prime, that is,
@@
Theorem 4.50.
for $\alpha=0$ or $\alpha\geq 3$, or if
then the equation (4.35) has no solutions in non-zero pairwise coprime integers $(x,y,z)$.
@@
Table 4.2: Values for $a,b,p,\alpha$ not covered by Theorem 4.50.
@@
and the values $ab$, $p$ and $\alpha$ are not those given in Table 4.3. Then the equation (4.35) has no solutions in non-zero pairwise coprime integers $(x,y,z)$.
@@
In 2014, Zhang and Luo [149] obtained the following result in the case $ab=2^{\alpha}3^{\beta}$ of equation (4.35).
@@
Suppose that $a$ and $b$ are coprime integers with $ab=2^{\alpha}3^{\beta}$ where $\alpha\geq 6$ and $\beta\geq 0$, and $p\geq 7$ is prime satisfying $p\nmid ab$. Then the equation (4.35) has no solutions in non-zero pairwise coprime integers $(x,y,z)$.
@@ 4.5 The general $a,b,c$
In this case, we are considering the equation (4.1) with general positive coefficients $a,b,c$. In this section, we will sometimes consider equations with $p$ or $r$ composite. If $p$ is even, we may also have $b<0$.
@@
Noubissie and Togbé [108, 109] studied equation (4.36) with $a=5^{\alpha}$, $\alpha\geq 1$, $b=64$, and $c=3$, and obtained the following result.
@@
has no non-zero primitive solutions with $xy\equiv 1$ (mod 2).
In 2006, Ivorra and Kraus [84] considered equation (4.36) with
@@
They provide some restrictions on $(l,n)$ and for $m\geq 1$, such that for $p$ larger than an explicit bound which depends on $n$, $l$ and $m$, the equation (4.36) has no non-trivial primitive integer solutions.
Ivorra and Kraus [84] also solved equation (4.36) for certain values of $(a,b,c)$ not of the form (4.37). Specifically, they obtain the following results.
@@
    For $p\geq 7$ and $(a,b,c)=(4,1,3)$, equation (4.36) only has the non-trivial primitive solutions $(x,y,z)=(1,-1,\pm 1)$.
@@
    For $p\geq 11$ and $(a,b,c)=(64,1,7)$, equation (4.36) only has the non-trivial primitive solutions $(x,y,z)=(1,-1,\pm 3)$.
In 2022, Chałupka et.al [37] studied equation (4.36) with $(a,b,c,p)=(3^{31},-4,7,34)$ and proved the following theorem.
@@
We next discuss equation (4.1) with $r=3$, that is,
@@
Noubissie and Togbé [108, 109] considered equation (4.38) with $a=2^{\alpha}$, $\alpha\geq 1$, $b=27$, $c=s^{\beta}$, $\beta\geq 1$, and $s\in\{7,13\}$, and proved the following results.
@@ 5 Equations of signature $(p,q,r)$
This section discusses equation (1.2) with exponents $p\neq q\neq r$. We will sometimes consider equations with some of the exponents being composite.
@@ 5.1 The case $a=c=1$
In this case, the equation (1.2) reduces to
@@
In 2014, Zhang and Luo [149] studied equation (5.1) for $p=2$, $r=3$ and $q$ even, that is,
@@
If $b=2^{\alpha}3^{\beta}31^{\gamma}$ and $q\nmid b$, then the equation (5.2) has no non-trivial solutions in pairwise coprime integers $x,y,z$. Also, if $\gamma=0$, then the same result holds for $q=7$.
Noubissie [107] studied the equation (5.1) with $p$ even, $q=2$ and $b=r=3$, that is
@@
The equation (5.3) with integer $p\geq 8$ has no non-trivial primitive integer solutions.
@@
The equation (5.1) with $b$ a positive odd integer, $p$ even, and $5\nmid h(-b)$ where $h(-b)$ denotes the class number of $\mathbb{Q}(\sqrt{-b})$ was considered by Zhang and Bai [147] who obtained the following partial result.
@@
In the same paper, Zhang and Bai [147] considered the equation (5.1) with $p=2$, $q$ even and $r=5$ and proved the following theorem.
@@
The authors of [149] also considered the equation (5.1) with $q=2p$ and $r=2$, that is
@@
Let $p\geq 7$ be prime, and let $b=4\cdot 3^{2\beta+1}$ with $\beta\geq 0$. Then the equation (5.4) has no non-trivial solutions in pairwise coprime integers $x,y,z$.
Noubissie [107] studied the equation (5.4) with $b=3^{2p-6}$, that is,

### v2

@@ 4.2 The case $a=c=1$
In this case, Eq. (4.1) is of the form
@@
Eq. (4.19) with $r=2$, $b=2^{\alpha}$ for $\alpha\geq 2$, and $p\geq 7$ prime was studied by Cohen [47, Section 15.3] who obtained the following result.
@@
Theorem 4.35 describes pairwise coprime solutions to (4.20) under the stated conditions, but this equation can have primitive solutions that are not pairwise coprime, e.g. $(x,y,z)=(2,1,3\cdot 2^{(p-3)/2})$ for $\alpha=p-3$, see [83]. In 2004, Ivorra [84, Chapter 1] proves that there are no other such examples, which completely solves Eq. (4.20) for $\alpha\geq 2$ and $p\geq 7$. Siksek also independently proved the same result in [131].
@@
If $1\leq\alpha\leq 4$ then for Eq. (4.20) with $p=5$, that is,
@@
Bennett, Vatsal and Yazdani [16] considered Eq. (4.19) with $r=3$ and $b$ being a prime power, that is,
$x^{p}+s^{\alpha}y^{p}=z^{3},$ (4.21)
@@
If $s$ and $p$ are primes such that $s\neq s_{1}^{3}\pm 3^{s_{2}}$ for any integers $s_{1}$ and $s_{2}$ with $s_{2}\neq 1$, $p>s^{2s}$ and $\alpha\geq 0$, then Eq. (4.21) has no solutions in coprime integers $x$, $y$ and $z$ with $|xy|>1$.
@@
Let $\alpha\geq 1$ be an integer,
$s\in\{5,11,13,23,29,31,41,43,47,53,59,61,67,71,79,83,97\}$
and let $p\geq 11$ be a prime not dividing $s^{2}-1$. Assume also that
[…]
(61,61),(67,73),(79,97),(97,13),(97,79)}.(61,61),(67,73),(79,97),(97,13),(97,79)\}.( 61 , 61 ) , ( 67 , 73 ) , ( 79 , 97 ) , ( 97 , 13 ) , ( 97 , 79 ) } .
Then Eq. (4.21) has no solutions in coprime integers $x$, $y$ and $z$ with $|xy|>1$.
So far in this section we considered the cases $r=2$ and $r=3$. For the next case $r=5$, we have the following result [137, Theorem 2.8].
@@
Azon [2] considered Eq. (4.19) where $r$ varies. Specifically, he considered Eq. (4.19) with $b=7$, $p=5$ and $r>41$ a prime and he proved that the equation has no non-trivial primitive integer solutions such that $10|z$.
@@ 4.3 The case $a=1$
In this instance, Eq. (4.1) reduces to
@@
We first consider the case $r=2$, which reduces Eq. (4.22) to
@@
Ivorra [83] studies Eq. (4.23) with $b=2^{\beta}$ for $10\leq\beta\leq p-$ 1, $c=2$ and $p\geq 7$ prime, that is,
@@
For prime $p\geq 7$ and integer $11\leq\beta\leq p-$ 1, Eq. (4.24) has no non-trivial primitive integer solutions.
Let prime $p\geq 7$ and $\alpha\geq 2$. The task of finding all primitive integer solutions to Eq. (4.23) with $b=2^{\alpha}$ and $c=3$, that is,
@@
is left as an exercise to the reader in [47, Section 15.8]. The authors of [154] prove that the equation has no solutions in coprime non-zero integers $x,y,z$ with $|xy|>1$.
Bennett and Skinner [15] obtained the following results concerning Eq. (4.23).
@@
$\alpha\geq\alpha_{0}$, $p>c$ and $(c,\alpha,p)\neq(11,3,13)$, then Eq. (4.23) has no solutions in non-zero pairwise coprime integers $(x,y,z)$ with $xy\neq\pm 1$.
@@
Suppose that $p\geq 11$ is prime and $\alpha$ is a non-negative integer. If $b\in\{5^{\alpha},11^{\alpha},13^{\alpha}\}$ and $p$ is not a divisor of $b$, then Eq. (4.23) with $c=2$ has no solution in non-zero pairwise coprime integers $(x,y,z)$.
Eq. (4.23) with $b$ a power of $2$ and $c$ a prime was studied by Zhang [151], who proved the following theorem.
@@
Bennett et al. [16] studied (4.26) with $b=s^{\alpha}$ and $c=3^{\beta}$ for prime $s$ and positive integers $\alpha,\beta$, that is,
@@
If $s$ and $p$ are primes such that $s\not\in\{5,3t^{3}\pm 1,9t^{3}\pm 1:t\in\mathbb{N}\}$, $p>s^{28s}$ and $\alpha,\beta$ are positive integers with $3\nmid\beta$, then Eq. (4.27) has no solutions in coprime integers $x$, $y$ and $z$ with $|xy|>1$.
@@
If $p\geq 7$ is prime, $s\in\{7,11,13\}$ with
$(s,p)\not\in\{(7,13),(13,7)\}$
and $\alpha,\beta$ are positive integers with $3\nmid\beta$, then Eq. (4.27) has no solutions in coprime integers $x$, $y$ and $z$ with $|xy|>1$.
Bennett et al. [16] also studied equation
@@
and $\alpha$ and $\beta$ are non-negative integers, then Eq. (4.28) has no solutions in coprime integers $x$, $y$ and $z$ with $|xy|>1$, unless $(|x|,|y|,\alpha,p,|s^{\beta}z^{3}|)=(2,1,1,7,125)$.
Krawciów [94] studied Eq. (4.26) with $b=s^{\alpha}$, that is,
@@
and solved it under some conditions on the parameters. To state the conditions, we must first introduce some notation. Let $s_{1},\ldots,s_{k}$ be odd primes, such that $3\in\{s_{1},\ldots,s_{k}\}$. Then let $P(s_{1},\ldots,s_{k})$ denote the set of primes $x$ satisfying any of the $2^{k+1}$ equations
$x^{\alpha}-{s_{1}}^{\alpha_{1}}\cdots{s_{k}}^{\alpha_{k}}y^{3}=\pm 1,\quad(1\leq\alpha_{i}\leq 2,\,\,i=1,\ldots,k),$
for some integers $\alpha>1$ and $y>1$. Theorem 4.34 implies that the set $P(s_{1},\ldots,s_{k})$ is always finite. For example, $P(3)=\{2,5\}$, see [94].
@@
Let $c=\prod_{i=1}^{k}s_{i}^{\gamma_{i}}$ be a prime factorization of a positive cube-free odd integer $c$ which is divisible by $3$, $\alpha$ be a positive integer, $p$ is a prime, and let $C=\mathrm{rad}(c)$. If $s$ is a prime such that $s\not\in P(s_{1},\dots,s_{k})$ and $s\neq\prod_{i=1}^{k}s_{i}^{\alpha_{i}}t^{3}\pm 1$ $(1\leq\alpha_{i}\leq 2,i=1,\dots,k)$ for any integer $t$, and if $p>(sC)^{10sC^{2}}$, then Eq. (4.29) has no non-trivial primitive integer solutions.
@@ 4.4 The case $c=1$
In this instance, Eq. (4.1) reduces to
@@
4.4.1 Eq. (4.30) with fixed $p$
Eq. (4.30) with $p=3$ was considered by Bennett and Dahmen which led to the following partial result [10, Theorem 13.2].
@@
has no solution in non-zero integers $x$, $y$ and $z$ (with $\gcd(x,y)=1$) and prime $r\equiv 1$ (mod 3) with $r\geq r_{0}$. If all solutions to (4.31) with $x$ and $y$ coprime integers have $\alpha_{2}\leq 4$, then there exists an effectively computable constant $r_{1}$ such that Eq. (4.32) has no solutions in odd integers $x$, $y$ and prime $r\geq r_{1}$.
@@
4.4.2 Eq. (4.30) with fixed $r$
In 2004, Bennett and Skinner [15] considered Eq. (4.30) with $r=2$, and $p\geq 11$ prime, that is,
@@
Table 4.2: Values for $a,b,p,\alpha$ not covered by Theorem 4.50.
@@
Theorem 4.50.
for $\alpha=0$ or $\alpha\geq 3$, or if
then Eq. (4.33) has no solutions in non-zero pairwise coprime integers $(x,y,z)$.
@@
and the values $ab$, $p$ and $\alpha$ are not those given in Table 4.3. Then Eq. (4.33) has no solutions in non-zero pairwise coprime integers $(x,y,z)$.
In 2014, Zhang and Luo [154] obtained the following result in the case $ab=2^{\alpha}3^{\beta}$ of Eq. (4.33).
@@
Suppose that $a$ and $b$ are coprime integers with $ab=2^{\alpha}3^{\beta}$ where $\alpha\geq 6$ and $\beta\geq 0$, and $p\geq 7$ is prime. Then Eq. (4.33) has no solutions in non-zero pairwise coprime integers $(x,y,z)$.
@@ 4.5 The general $a,b,c$
In this case, we are considering Eq. (4.1) with general positive coefficients $a,b,c$. In this section, we will sometimes consider equations with $p$ or $r$ composite. If $p$ is even, we may also have $b<0$.
@@
Noubissie and Togbé [111, 112] studied Eq. (4.34) with $a=5^{\alpha}$, $\alpha\geq 1$, $b=64$, and $c=3$, and obtained the following result.
@@
has no non-zero primitive integer solutions with $xy\equiv 1$ (mod 2).
In 2006, Ivorra and Kraus [86] considered Eq. (4.34) with
@@
They provide some restrictions on $(l,n)$ and for $m\geq 1$, such that for $p$ larger than an explicit bound which depends on $l$, $m$ and $n$, Eq. (4.34) has no non-trivial primitive integer solutions.
Ivorra and Kraus [86] also solved Eq. (4.34) for certain values of $(a,b,c)$ not of the form (4.35). Specifically, they obtain the following results.
@@
    For $p\geq 7$ and $(a,b,c)=(4,1,3)$, Eq. (4.34) only has the non-trivial primitive integer solutions $(x,y,z)=(1,-1,\pm 1)$.
@@
    For $p\geq 11$ and $(a,b,c)=(64,1,7)$, Eq. (4.34) only has the non-trivial primitive integer solutions $(x,y,z)=(1,-1,\pm 3)$.
In 2022, Chałupka et al. [38] studied Eq. (4.34) with $(a,b,c,p)=(3^{31},-4,7,34)$ and proved the following theorem.
@@
We next discuss Eq. (4.1) with $r=3$, that is,
@@
Noubissie and Togbé [111, 112] considered Eq. (4.36) with $a=2^{\alpha}$, $\alpha\geq 1$, $b=27$, $c=s^{\beta}$, $\beta\geq 1$, and $s\in\{7,13\}$, and proved the following results.
@@
We next discuss Eq. (4.1) with $r=5$, that is,
$ax^{p}+by^{p}=cz^{5}.$ (4.37)
Azon [2] considered Eq. (4.37) with $a=7$, $b=1$, $c=3$ and $p>71$ a prime and he proved that this equation has no non-trivial primitive integer solutions such that $10|xy$. In the same paper, he also proved the following result.
Theorem 4.58.
  • (a)
    Suppose that $p>71$ is a prime. Then, for any $\alpha\in\{1,2,3,4\}$ and $\beta\in\{3,4\}$, there are no non-trivial primitive integer solutions to the equation
    $7x^{p}+2^{i}5^{j}y^{p}=3z^{5}.$
  • (b)
    Suppose that $p>41$ is a prime. Then, for any $\alpha\in\{1,2,3,4\}$ and $\beta\in\{2,3,4\}$, there are no non-trivial primitive integer solutions to the equation
    $x^{5}+7y^{5}=2^{\alpha}5^{\beta}z^{p}.$
@@ 5 Equations of signature $(p,q,r)$
This section discusses Eq. (1.1) with exponents $p\neq q\neq r$. We will sometimes consider equations with some of the exponents being composite.
@@ 5.1 The case $a=c=1$
In this case, Eq. (1.1) reduces to
@@
In 2014, Zhang and Luo [154] studied Eq. (5.1) for $p=2$, $r=3$ and $q$ even, that is,
@@
If $b=2^{\alpha}3^{\beta}31^{\gamma}$ and $q\nmid b$, then Eq. (5.2) has no non-trivial solutions in pairwise coprime integers $x,y,z$. Also, if $\gamma=0$, then the same result holds for $q=7$.
Noubissie [110] studied Eq. (5.1) with $p$ even, $q=2$ and $b=r=3$, that is,
@@
Eq. (5.3) with integer $p\geq 8$ has no non-trivial primitive integer solutions.
@@
Eq. (5.1) with $b$ a positive odd integer, $p$ even, and $5\nmid h(-b)$ where $h(-b)$ denotes the class number of $\mathbb{Q}(\sqrt{-b})$ was considered by Zhang and Bai [152] who obtained the following partial result.
@@
In the same paper, Zhang and Bai [152] considered Eq. (5.1) with $p=2$, $q$ even and $r=5$ and proved the following theorem.
@@
The authors of [154] also considered Eq. (5.1) with $q=2p$ and $r=2$, that is,
@@
Let $p\geq 7$ be prime, and let $b=4\cdot 3^{2\beta+1}$ with $\beta\geq 0$. Then Eq. (5.4) has no non-trivial solutions in pairwise coprime integers $x,y,z$.
Noubissie [110] studied Eq. (5.4) with $b=3^{2p-6}$, that is,